OpenStudy (anonymous):

what is the derivative of (lnx^2)/(x+1)^(1/2)

OpenStudy (amistre64):

\[D_x(\frac{\ln(x)}{(x+1)^2})\]

OpenStudy (amistre64):

this thing? and whered my fraction bar end up :)

OpenStudy (amistre64):

use quotient rule; or bring the botom up with a negative exponent and use product rule

OpenStudy (amistre64):

murmuring..... ^(1/2) lol

OpenStudy (amistre64):

bt'-b't f'(x) = ----- b^2

OpenStudy (amistre64):

b = (x+1)^(1/2) ; b' = 1/[2(x+1)^(1/2)] t = ln(x) ; t' = 1/x

OpenStudy (amistre64):

b^2 = (x+1) :) that one was easy lol

OpenStudy (amistre64):

\[\frac{\frac{\sqrt{x+1}}{x} - \frac{\ln(x)}{2 \sqrt{x+1}}}{(x+1)}\] simplify as needed

OpenStudy (amistre64):

2x + 2 - x ln(x) --------------- 2x(x+1) sqrt(x+1)

OpenStudy (anonymous):

amistre64, In the following equation, Ln is spelled Log. \[\int\limits \frac{\frac{\sqrt{1+x}}{x}-\frac{\text{Log}[x]}{2 \sqrt{1+x}}}{1+x} \, dx = \frac{\text{Log}[x]}{\sqrt{1+x}} \] It looks to me like the Ln of x was raised to the second power in the original problem statement above. Perhaps by eyes are failing me in my old age.